# Non-Rectangular Average-Reward Robust MDPs: Optimal Policies and Their Transient Values

### 03 Mar 2026 v2

@@ 1 Introduction
This paper addresses these challenges in non-rectangular average-reward robust MDPs. We impose no structural assumptions on the ambiguity set beyond weak communication, which guarantees that, in the non-robust case, the optimal average reward is well defined and independent of the initial state. Recognizing that Markovian policies may fail to be optimal under non-rectangular ambiguity, we adopt the standard formulation in Grand-Clement et al., [7]: the controller may employ general history-dependent policies, while the adversary is stationary and commits to a single transition kernel $p\in\mathcal{P}$ for the entire horizon. This non-rectangular model encompasses a broad class of statistically natural ambiguity models, including those induced by joint confidence regions, maximum-likelihood estimators, or nonlinear factor models, while lying outside the classical robust dynamic programming framework that relies on rectangularity to ensure stagewise decomposability.
@@ 1.1 Review of Relevant Literature
Rectangular robust MDPs: The discounted-reward theory for robust MDPs under rectangular ambiguity sets is now well developed: Rectangularity restores a dynamic programming principle and leads to Bellman-type optimality equations [8, 26, 27, 22, 6, 23]. By contrast, the average-reward setting has received comparatively less attention and has only recently begun to be systematically developed. For example, Wang et al., 2023b [24] study the average-reward robust problem under SA-rectangularity together with uniform unichain assumptions. Grand-Clement et al., [7] investigate Blackwell optimality, proving that $\epsilon$-Blackwell optimal policies exist under SA-rectangularity, and highlighting important pathologies beyond this regime: in S-rectangular robust MDPs, average-reward optimal policies may fail to exist. More recently, Wang and Si, [21] develops dynamic programming and Bellman equations for average-reward formulations under broad conditions within the rectangular framework. Taken together, these works highlight both the analytical strength of rectangularity and the fragility of average-reward optimality, thereby motivating our study of non-rectangular ambiguity under a stationary committing adversary in the average-reward setting.
@@ 2.1 Dynamic Programming for Classical Average-Reward MDPs
Since it is assumed that $\left\{{\delta_{a}:a\in A}\right\}\subseteq\mathcal{Q}$, the dynamic programming principle for average-reward MDP (see, e.g., Puterman, [16, Theorem 9.1.4 and 9.1.7]) implies that
@@
for all $s\in S$. A solution to (2.2) always exists when the state and action spaces are finite [16, Theorem 9.1.4 and Proposition 9.1.1]. Moreover, in this setting, $\alpha_{p}$ is unique in the sense that if $(\alpha^{\prime},v^{\prime})$ also solves (2.2), then $\alpha_{p}=\alpha^{\prime}$.
@@ 3.1 Online RL Algorithms and Expected Regret
where $\alpha_{p}$ is the unique solution to (2.2). Then the expected regret of a policy $\pi$ is given by $E_{\mu}^{\pi,p}R_{r,p}(T)$. Because $S$ and AA are finite, any history-dependent policy has expected regret $E_{\mu}^{\pi,p}R_{r,p}(T)=O(T)$ as $T\to\infty$.
@@ 3.3 Weak Communication and the Existence of Online RL Policies
First, note that for $k=1,2$, under kernel $p^{(k)}$ and starting from state $i$, taking action $a_{k}$ always leads to $e_{1}$ as the next state. Since $e_{1}$ is an absorbing state with reward 11, the long-run average reward under this action is 11. Consequently, $\alpha_{p}(i)=1$ for all $p\in\mathcal{P}$.
@@
$P_{\mu}^{\pi(T),p}\left(\frac{R_{r,p}(T)}{f_{\mu,r,p}T^{\eta}}\leq 1\right)\;\geq\;1-\frac{1}{T^{\beta}}.$
@@
for any $\mu\in\mathcal{P}(S)$, $r:S\times A\to[0,1]$, and weakly communicating pp. This inequality bounds the regret of $\pi(T)$ over any horizon of length $T$.
@@
Suppose $\left\{{\delta_{a}:a\in A}\right\}\subseteq\mathcal{Q}$ and $\mathcal{P}$ is weakly communicating. Then Zhang and Xie, [28, Theorem 1] shows that the UCB-AVG algorithm, identified as $\pi_{\text{ZX}}^{\delta}\in\Pi_{\mathrm{H}}(\mathcal{Q})$, which only uses deterministic actions, achieves
@@
for all initial distributions $\mu$ and all $T\geq 1$. Here $C$ is a constant that only depends on $|S|$ and $|A|$, and $v_{p}^{*}:S\rightarrow\mathbb{R}$ is one of the solutions to (3.5), which depends on $r$ and pp.
@@ 4.1 Transient Values and Their Upper Bounds
To show the final statement, we construct a simple robust MDP instance with $S=\left\{{s}\right\}$ and $A=\left\{{a_{0},a_{1}}\right\}$. In this setting, there is only one possible transition kernel $p$, defined by $p(s|s,a_{i})=1$ for $i=0,1$. Therefore, the only possible ambiguity set is $\mathcal{P}=\left\{{p}\right\}$. We let $\mathcal{Q}=\mathcal{P}(\mathcal{A})$ or $\left\{{\delta_{a_{0}},\delta_{a_{1}}}\right\}$, and consider a reward function $r(s,a_{i})=r(a_{i})=i$ for $i=0,1$. It is immediate that the optimal average reward is 11, achieved by the policy that always plays action $a_{1}$.
@@ 4.2 Transient Value Lower Bounds for RL Policies
In particular, a typical minimax-optimal RL policy achieves $\widetilde{O}(\sqrt{T})$ [9, 28] regret for all weakly communicating $p\in\mathcal{P}$, with a common growth rate but instance-dependent constants. In this case, one may take $w_{p}(T)=\widetilde{\Omega}(T^{-1/2})$ for all such pp. For instance, applying Proposition 4.2 to the RL policy $\pi_{\mathrm{ZX}}$ in Example 2, we obtain that under the assumptions of Corollary 1.1, for any weight function $w(T)=O(T^{-1/2-\epsilon})$, we have $\mathrm{TV}_{w}(\mu,\pi_{\mathrm{ZX}})=0$ for all $\mu\in\mathcal{P}(\mathcal{S})$.
@@ 5.1.1 Proof of Theorem 2
Next, we prove the second statement of Theorem 2. We first present Lemma 2, which states an important property of the product Dirichlet prior in Definition 5. This property underlies our analysis by turning the mixture likelihood ratio process $\Lambda_{n}$ into a likelihood ratio process for testing against a simple alternative $P$. The proof of Lemma 2 is deferred to A.2.
@@ 5.3.1 Proof of Theorem 3
We first note that, with unichain $P_{0}$, the existence and up-to-shift uniqueness of $v^{*}$ is standard for finite-state unichain average-reward models; see, e.g., [16, Chs. 8–9].
@@ Acknowledgments
Shengbo Wang would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Bridging Stochastic Control And Reinforcement Learning, where work on this paper was undertaken. This work was supported by EPSRC grant EP/V521929/1.

### 10 Mar 2026 v3

@@ 1 Introduction
This paper addresses these challenges in non-rectangular average-reward robust MDPs. We impose no structural assumptions on the ambiguity set beyond weak communication, which guarantees that, in the non-robust case, the optimal average reward is well defined and independent of the initial state. Recognizing that Markovian policies may fail to be optimal under non-rectangular ambiguity, we adopt the standard formulation in Grand-Clément et al., [7]: the controller may employ general history-dependent policies, while the adversary is stationary and commits to a single transition kernel $p\in\mathcal{P}$ for the entire horizon. This non-rectangular model encompasses a broad class of statistically natural ambiguity models, including those induced by joint confidence regions, maximum-likelihood estimators, or nonlinear factor models, while lying outside the classical robust dynamic programming framework that relies on rectangularity to ensure stagewise decomposability.
@@ 1.1 Review of Relevant Literature
Rectangular robust MDPs: The discounted-reward theory for robust MDPs under rectangular ambiguity sets is now well developed: Rectangularity restores a dynamic programming principle and leads to Bellman-type optimality equations [10, 28, 29, 24, 8, 25]. By contrast, the average-reward setting has received comparatively less attention and has only recently begun to be systematically developed. For example, Wang et al., 2023b [26] study the average-reward robust problem under SA-rectangularity together with uniform unichain assumptions. Grand-Clément et al., [7] investigate Blackwell optimality, proving that $\epsilon$-Blackwell optimal policies exist under SA-rectangularity, and highlighting important pathologies beyond this regime: in S-rectangular robust MDPs, average-reward optimal policies may fail to exist. Relatedly, Grand-Clément and Vieille, [9] show that for $s$-rectangular robust MDPs, Blackwell $\epsilon$-optimal policies may fail to be Markovian, while full history dependence is unavoidable in general. More recently, Wang and Si, [23] develops dynamic programming and Bellman equations for average-reward formulations under broad conditions within the rectangular framework. Taken together, these works highlight both the analytical strength of rectangularity and the fragility of average-reward optimality, thereby motivating our study of non-rectangular ambiguity under a stationary committing adversary in the average-reward setting.
@@ 2.1 Dynamic Programming for Classical Average-Reward MDPs
Since it is assumed that $\left\{{\delta_{a}:a\in A}\right\}\subseteq\mathcal{Q}$, the dynamic programming principle for average-reward MDP (see, e.g., Puterman, [18, Theorem 9.1.4 and 9.1.7]) implies that
@@
for all $s\in S$. A solution to (2.2) always exists when the state and action spaces are finite [18, Theorem 9.1.4 and Proposition 9.1.1]. Moreover, in this setting, $\alpha_{p}$ is unique in the sense that if $(\alpha^{\prime},v^{\prime})$ also solves (2.2), then $\alpha_{p}=\alpha^{\prime}$.
@@ 3.1 Online RL Algorithms and Expected Regret
where $\alpha_{p}$ is the unique solution to (2.2). Then the expected regret of a policy $\pi$ is given by $E_{\mu}^{\pi,p}R_{r,p}(T)$. Because $S$ and $A$ are finite, any history-dependent policy has expected regret $E_{\mu}^{\pi,p}R_{r,p}(T)=O(T)$ as $T\to\infty$.
@@ 3.3 Weak Communication and the Existence of Online RL Policies
First, note that for $k=1,2$, under kernel $p^{(k)}$ and starting from state $i$, taking action $a_{k}$ always leads to $e_{1}$ as the next state. Since $e_{1}$ is an absorbing state with reward $1$, the long-run average reward under this action is $1$. Consequently, $\alpha_{p}(i)=1$ for all $p\in\mathcal{P}$.
@@
$P_{\mu}^{\pi(T),p}\left(\frac{R_{r,p}(T)}{f_{\mu,r,p}T^{\eta}}\leq 1\right)\geq 1-\frac{1}{T^{\beta}}.$
@@
for any $\mu\in\mathcal{P}(S)$, $r:S\times A\to[0,1]$, and weakly communicating $p$. This inequality bounds the regret of $\pi(T)$ over any horizon of length $T$.
@@
Suppose $\left\{{\delta_{a}:a\in A}\right\}\subseteq\mathcal{Q}$ and $\mathcal{P}$ is weakly communicating. Then Zhang and Xie, [30, Theorem 1] shows that the UCB-AVG algorithm, identified as $\pi_{\text{ZX}}^{\delta}\in\Pi_{\mathrm{H}}(\mathcal{Q})$, which only uses deterministic actions, achieves
@@
for all initial distributions $\mu$ and all $T\geq 1$. Here $C$ is a constant that only depends on $|S|$ and $|A|$, and $v_{p}^{*}:S\rightarrow\mathbb{R}$ is one of the solutions to (3.5), which depends on $r$ and $p$.
@@ 4.1 Transient Values and Their Upper Bounds
We note that Assumption 1 serves as the minimal condition that makes the definition (4.1) meaningful when $w(T)$ is not $O(1/T)$. Indeed, if there is no $p\in\mathcal{P}$ such that $\alpha^{*}(\mu)=\sum_{s\in S}\mu(s)\alpha_{p}(s)$, then for any ${\pi_{\mathrm{RL}}}\in{\Pi_{\mathrm{RL}}}$,
$E_{\mu}^{{\pi_{\mathrm{RL}}},p}\sum_{t=0}^{T-1}[r(X_{t},A_{t})-\alpha^{*}(\mu)]=E_{\mu}^{{\pi_{\mathrm{RL}}},p}\sum_{t=0}^{T-1}[r(X_{t},A_{t})-\alpha_{p}(X_{0})]+T\left(\sum_{s\in S}\mu(s)\alpha_{p}(s)-\alpha^{*}(\mu)\right)=\Omega(T)$
as $T\rightarrow\infty$.
On the other hand, even if $\mathcal{P}$ is compact and convex, a worst-case kernel need not exist [7, Proposition 3.2]. However, existence can be guaranteed when $\mathcal{P}$ is weakly communicating. We summarize this fact in the following lemma as a side result. To streamline the discussion of transient values, we defer the proof of Lemma 1 to Appendix A.1.
Lemma 1 (Existence of Worst-Case Kernels).
Assume that $\mathcal{P}$ is compact and weakly communicating. Then Assumption 1 holds.
With a well-defined finite transient value under Assumption 1, we establish some important properties.
@@
To show the final statement, we construct a simple robust MDP instance with $S=\left\{{s}\right\}$ and $A=\left\{{a_{0},a_{1}}\right\}$. In this setting, there is only one possible transition kernel $p$, defined by $p(s|s,a_{i})=1$ for $i=0,1$. Therefore, the only possible ambiguity set is $\mathcal{P}=\left\{{p}\right\}$. We let $\mathcal{Q}=\mathcal{P}(\mathcal{A})$ or $\left\{{\delta_{a_{0}},\delta_{a_{1}}}\right\}$, and consider a reward function $r(s,a_{i})=r(a_{i})=i$ for $i=0,1$. It is immediate that the optimal average reward is $1$, achieved by the policy that always plays action $a_{1}$.
@@ 4.2 Transient Value Lower Bounds for RL Policies
In particular, a typical minimax-optimal RL policy achieves $\widetilde{O}(\sqrt{T})$ [11, 30] regret for all weakly communicating $p\in\mathcal{P}$, with a common growth rate but instance-dependent constants. In this case, one may take $w_{p}(T)=\widetilde{\Omega}(T^{-1/2})$ for all such $p$. For instance, applying Proposition 4.2 to the RL policy $\pi_{\mathrm{ZX}}$ in Example 2, we obtain that under the assumptions of Corollary 1.1, for any weight function $w(T)=O(T^{-1/2-\epsilon})$, we have $\mathrm{TV}_{w}(\mu,\pi_{\mathrm{ZX}})=0$ for all $\mu\in\mathcal{P}(\mathcal{S})$.
@@ 5.1.1 Proof of Theorem 2
Next, we prove the second statement of Theorem 2. We first present Lemma 3, which states an important property of the product Dirichlet prior in Definition 5. This property underlies our analysis by turning the mixture likelihood ratio process $\Lambda_{n}$ into a likelihood ratio process for testing against a simple alternative $P$. The proof of Lemma 3 is deferred to A.3.
@@ 5.3.1 Proof of Theorem 3
We first note that, with unichain $P_{0}$, the existence and up-to-shift uniqueness of $v^{*}$ is standard for finite-state unichain average-reward models; see, e.g., [18, Chs. 8–9].
@@ Acknowledgments
We thank Julien Grand-Clément for valuable input and comments. Nian Si’s research is supported in part by the Early Career Scheme [Grant 26210125] from the Hong Kong Research Grants Council. Shengbo Wang would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Bridging Stochastic Control And Reinforcement Learning, where work on this paper was undertaken. This work was supported by EPSRC grant EP/V521929/1.
@@ A.1 Proof of Lemma 1
Proof.
Recall that under weak communication, $\alpha_{p}(s)=\alpha_{p}$ for all $s\in S$, where $(\alpha_{p},v_{p})$ solves (3.5) and $\alpha_{p}$ is unique. Moreover, by Corollary 1.1, $\alpha^{*}(\mu)=\alpha^{*}=\inf_{p\in\mathcal{P}}\alpha_{p},$ for all $\mu\in\mathcal{P}(\mathcal{S}).$ Hence, since $\mathcal{P}$ is compact, it suffices to prove that $p\rightarrow\alpha_{p}$ is continuous on $\mathcal{P}$.
For $p,q\in\mathcal{P}$, define
$d(p,q):=\max_{s\in S,a\in A}d_{\mathrm{TV}}\bigl(p(\cdot|s,a),q(\cdot|s,a)\bigr),$
and, for $v:S\to\mathbb{R}$,
$\mathcal{B}_{p}[v](s):=\max_{a\in A}\left\{{r(s,a)+\sum_{s^{\prime}\in S}p(s^{\prime}|s,a)v(s^{\prime})}\right\}.$
Since
$\left|\sum_{s^{\prime}\in S}[q(s^{\prime}|s,a)-p(s^{\prime}|s,a)]v(s^{\prime})\right|\leq d(p,q)\left|v\right|_{\mathrm{span}},$
we have, for every $s\in S$,
$\mathcal{B}_{q}[v_{p}](s)\leq\mathcal{B}_{p}[v_{p}](s)+d(p,q)\left|v_{p}\right|_{\mathrm{span}}=v_{p}(s)+\beta_{p,q}.$
where $\beta_{p,q}:=\alpha_{p}+d(p,q)\left|v_{p}\right|_{\mathrm{span}}$ and we use (3.5) that $\mathcal{B}_{p}[v_{p}]-\alpha_{p}=v_{p}(s)$. Thus $v_{p}$ satisfies the Bellman inequality for the controlled kernel $q$.
By the comparison theorem for average-reward MDPs [18, Theorem 9.1.2],
$\alpha_{q}\leq\beta_{p,q}=\alpha_{p}+d(p,q)\left|v_{p}\right|_{\mathrm{span}}.$
Exchanging the roles of $p$ and $q$ gives
$\alpha_{p}\leq\alpha_{q}+d(p,q)\left|v_{q}\right|_{\mathrm{span}}.$
Hence
$|\alpha_{q}-\alpha_{p}|\leq d(p,q)\max\{\left|v_{p}\right|_{\mathrm{span}},\left|v_{q}\right|_{\mathrm{span}}\}.$ (A.1)
It remains to show that $\left|v_{q}\right|_{\mathrm{span}}$ is locally bounded around each $p\in\mathcal{P}$. Fix $p\in\mathcal{P}$, and let $C_{p}$ be the communicating class in the weak-communication decomposition of $p$. Write $\tau_{B}$ for the first hitting time of a set $B\subseteq S$.
First, we control the entrance time into $C_{p}$ under arbitrary stationary deterministic policies $\pi\in\Pi_{\mathrm{SD}}$. Let $Q_{q,\pi}$ be the restriction of the transition matrix under $(q,\pi)$ to $S\setminus C_{p}$. Since every state in $S\setminus C_{p}$ is transient under every stationary policy for $p$, the matrix $I-Q_{p,\pi}$ is invertible, and by a first transition argument
$(I-Q_{p,\pi})^{-1}e](s)=E_{s}^{\pi,p}\tau_{C_{p}}$
for all $s\in S\setminus C_{p}$, where $e$ is the the vector of all ones. As $q\rightarrow Q_{q,\pi}$ is continuous and $\Pi_{\mathrm{SD}}$ is finite, there exist $\delta_{p}^{(1)}>0$ and $M_{p}<\infty$ such that
$\sup_{q:d(p,q)\leq\delta_{p}^{(1)}}\max_{\pi\in\Pi_{\mathrm{SD}}}\max_{s\in S\setminus C_{p}}E_{s}^{\pi,q}\tau_{C_{p}}\leq M_{p}.$ (A.2)
Next, we control the hitting time of any prescribed state in $C_{p}$ under a suitable stationary policy. Fix $s^{\prime}\in C_{p}$, and consider the directed graph on $S$ with an edge $x\to y$ whenever $p(y|x,a)>0$ for some $a\in A$. Every state $x\in S$ has a directed path to $s^{\prime}$: this is clear for $x\in C_{p}$; if $x\notin C_{p}$, then it is transient and hence some $y\in C_{p}$ is reachable from $x$. But $y$ can reach any $s\in C_{p}$ by the definition of $C_{p}$. So, $x$ can reach $s^{\prime}$ through $y$.
Let $\ell_{s^{\prime}}(x)$ be the graph distance from $x$ to $s^{\prime}$. For each $x\neq s^{\prime}$, choose a successor $y_{x,s^{\prime}}$ and an action $a_{x,s^{\prime}}\in A$ such that
$\ell_{s^{\prime}}(y_{x,s^{\prime}})=\ell_{s^{\prime}}(x)-1,\qquad p(y_{x,s^{\prime}}|x,a_{x,s^{\prime}})>0.$
Define a stationary deterministic policy $\pi^{s^{\prime}}\in\Pi_{\mathrm{SD}}$ by $\pi^{s^{\prime}}(x)=a_{x,s^{\prime}}$ for $x\neq s^{\prime}$ and take arbitrary action at $s^{\prime}$. Under $\pi^{s^{\prime}}$, from every $x\in S$ there is a path to $s^{\prime}$ of length at most
$m_{p}(s^{\prime}):=\max_{x\in S}\ell_{s^{\prime}}(x)<\infty$
and with strictly positive probability under $(\pi^{s^{\prime}},p)$. Since there are only finitely many such designated paths, their probabilities admit a strictly positive minimum:
$\eta_{p}(s^{\prime}):=\min_{x\in S}P_{x}^{\pi^{s^{\prime}},p}(\tau_{s^{\prime}}\leq m_{p}(s^{\prime}))>0.$
By continuity of hitting probabilities in $q$, there exists $\delta_{p}^{(2)}(s^{\prime})>0$ such that, whenever $d(p,q)\leq\delta_{p}^{(2)}(s^{\prime})$,
$P_{x}^{\pi^{s^{\prime}},q}(\tau_{s^{\prime}}\leq m_{p}(s^{\prime}))\geq\eta_{p}(s^{\prime})/2,$
for all $x\in S.$ By the Markov property, this implies the geometric tail bound
$P_{x}^{\pi^{s^{\prime}},q}(\tau_{s^{\prime}}>km_{p}(s^{\prime}))\leq(1-\eta_{p}(s^{\prime})/2)^{k},\qquad k\geq 0.$
Therefore, let $\delta^{(2)}_{p}:=\min_{s^{\prime}\in C_{p}}\delta_{p}^{(2)}(s^{\prime})$
$\sup_{q:d(p,q)\leq\delta_{p}^{(2)}}\max_{s\in S,s^{\prime}\in C_{p}}E_{s}^{\pi^{s^{\prime}},q}\tau_{s^{\prime}}\leq\max_{s^{\prime}\in C_{p}}\frac{2m_{p}(s^{\prime})}{\eta_{p}(s^{\prime})}=:L_{p}.$ (A.3)
Now fix $q\in\mathcal{P}$ with
$d(p,q)\leq\delta_{p}:=\min\{\delta_{p}^{(1)},\delta_{p}^{(2)}\},$
and choose
$\bar{s}=\bar{s}(q)\in\arg\max_{s\in C_{p}}v_{q}(s).$
Let $\pi_{q}^{*}\in\Pi_{\mathrm{SD}}$ be a stationary deterministic policy attaining the maximum in the Bellman equation for $q$. Since $0\leq r\leq 1$, we also have $0\leq\alpha_{q}\leq 1$, hence $1\leq r-\alpha_{q}\leq 1$. Moreover the Bellman equation implies that
$G_{n}:=v_{q}(X_{n})+\sum_{t=0}^{n-1}[r(X_{t},\pi_{q}^{*}(X_{t}))-\alpha_{q}]$
is a $P_{s}^{\pi^{*}_{q},q}$ martingale.
Since $\delta_{p}\leq\delta_{p}^{(1)}$, $P_{s}^{\pi^{*}_{q},q}(\tau_{C_{p}}<\infty)=1$. Therefore, stopping at $\tau_{C_{p}}\wedge n$ and letting $n\uparrow\infty$, we obtain for every $s\in S$,
$v_{q}(s)=E_{s}^{\pi_{q}^{*},q}\!\left[\sum_{t=0}^{\tau_{C_{p}}-1}\bigl(r(X_{t},\pi_{q}^{*}(X_{t}))-\alpha_{q}\bigr)+v_{q}(X_{\tau_{C_{p}}})\right]\leq v_{q}(\bar{s})+E_{s}^{\pi_{q}^{*},q}\tau_{C_{p}},$
because $X_{\tau_{C_{p}}}\in C_{p}$ and $v_{q}(X_{\tau_{C_{p}}})\leq v_{q}(\bar{s})$. Therefore, by (A.2),
$\max_{s\in S}v_{q}(s)\leq v_{q}(\bar{s})+M_{p}.$ (A.4)
On the other hand, for any $s\in S$, apply the same arguments to the Bellman inequality
$v_{q}(s)\geq r(s,\pi^{\bar{s}}(s))-\alpha_{q}+\sum_{s^{\prime}\in S}q(s^{\prime}|s,\pi^{\bar{s}}(s))v_{q}(s^{\prime})$
and the stopping time $\tau_{\bar{s}}\wedge n$ and letting $n\uparrow\infty$ yields
$v_{q}(s)\geq E_{s}^{\pi^{\bar{s}},q}\!\left[\sum_{t=0}^{\tau_{\bar{s}}-1}\bigl(r(X_{t},\pi^{\bar{s}}(X_{t}))-\alpha_{q}\bigr)+v_{q}(\bar{s})\right]\geq v_{q}(\bar{s})-E_{s}^{\pi^{\bar{s}},q}\tau_{\bar{s}}.$
Hence, by (A.3),
$v_{q}(\bar{s})-\min_{s\in S}v_{q}(s)\leq\max_{s\in S}E_{s}^{\pi^{\bar{s}},q}\tau_{\bar{s}}\leq L_{p}.$ (A.5)
Combining (A.4) and (A.5), we conclude that
$\sup_{q:d(p,q)\leq\delta_{p}}\left|v_{q}\right|_{\mathrm{span}}\leq M_{p}+L_{p}<\infty.$
Returning to (A.1), for $d(p,q)\leq\delta_{p}$,
$|\alpha_{q}-\alpha_{p}|\leq d(p,q)\bigl(\left|v_{p}\right|_{\mathrm{span}}+M_{p}+L_{p}\bigr).$
Thus $q\to p$ implies $\alpha_{q}\to\alpha_{p}$. Therefore $p\rightarrow\alpha_{p}$ is continuous on $\mathcal{P}$. Since $\mathcal{P}$ is compact, there exists $p^{*}\in\mathcal{P}$ such that $\alpha_{p^{*}}=\min_{p\in\mathcal{P}}\alpha_{p}=\alpha^{*}.$ Therefore, Assumption 1 holds. ∎
A.2 Proof of Lemma 2